(5 pts) Let X denote the annual salary of scientists in the United States, measured in thousands of dollars. Suppose that EX] 52.5 and that the standard deviation of X is 12.5. Let Y denote the annual salary of scientists in the United States measured in dollars. What is EY and what is the standard deviation of Y?
Answer #1

here as we can see that Y=1000*X

therefore E(Y)=1000*E(X)=52.5*1000 =52500

and SD(Y)=SD(1000*X)=1000*SD(X) =1000*12.5=12500